3 years ago, Rica was thrice as old as Dandy. 3 years hence, she will be twice as old. Find their present ages.
Answer by mananth(16946)


3 years ago
Let Dandy be x years
Rica will be 3x
..
3 years later
Dandy will be x+6
Rica will be 3x+3
..
3x+6=2(x+6)
3x+6 =2x+12
3x-2x=12-6
x=6 Dandy's age 3 years ago. so she will be 9 years now.
Rica will be 3x+3 = 3*6+3= 21 years